# Bayesian nonparametric modeling of multivariate count data with an unknown number of traits

### v1

@@ 1 Introduction
Feature allocation models gained increasing popularity within the Bayesian nonparametric community since the definition of the Indian buffet process (ibp) by Griffiths and Ghahramani (2006, 2011). In these models, each subject (or observation) is characterized by the presence of a set of latent characteristics, referred to as features. A key development was due to Thibaux and Jordan (2007), who determined the de Finetti measure of the ibp. These seminal contributions paved the way for a rigorous theoretical foundation of exchangeable feature allocations (Broderick et al., 2013). Relevant extensions are the three-parameter ibp (Teh and Gorur, 2009), scaled processes (Masoero et al., 2022; Camerlenghi et al., 2024), and product-form feature models defined by Battiston et al. (2018) and studied by Ghilotti et al. (2025). See also Beraha et al. (2025) for a perspective based on point processes theory and predictive characterizations for feature models. Trait allocation models (James, 2017; Campbell et al., 2018) are an important generalization of feature allocation models, where each subject’s characteristics, referred to as traits, are associated with non-negative measurements rather than binary features. Notable examples of trait allocation models are described in Broderick et al. (2015); Heaukulani and Roy (2016); Beraha and Favaro (2025). Applications of feature and trait allocation models encompass several fields, including genomics (Beraha and Favaro, 2025; Masoero et al., 2022), Bayesian factor analysis (Ayed and Caron, 2021), ecology (Stolf and Dunson, 2025; Ghilotti et al., 2025), gene expression modeling (Knowles and Ghahramani, 2011), document classification and topic modeling (Zhou et al., 2016; Williamson et al., 2010), image segmentation and object recognition (Griffiths and Ghahramani, 2011; Broderick et al., 2015).
@@
Despite their popularity, most existing contributions on both feature and trait allocation models assume exchangeability among subjects, which implies a form of homogeneity across observations. In many applied areas, this assumption is restrictive, as data come from multiple related studies and are intrinsically heterogeneous. In such cases, the notion of partial exchangeability is more appropriate. Under this assumption, observations are divided into groups that are exchangeable within, but not across, the different groups. Partial exchangeability has a long history to handle data from dependent populations of species: refer e.g. to Lijoi et al. (2014); Griffin and Leisen (2017); Camerlenghi et al. (2019); Colombi et al. (2025); Franzolini et al. (2025) for specific structures, and to Quintana et al. (2022) for a complete review. Nonetheless, only a few and recent contributions are available for trait and feature allocation models. For instance, Masoero et al. (2018); Beraha and Favaro (2025); James et al. (2024) introduced hierarchical constructions to induce dependence among groups, while Shen et al. (2024) proposed a bivariate beta process.
@@
This paper introduces a novel general and tractable class of Bayesian nonparametric priors suitable for modeling partially exchangeable trait allocations. Our proposal relies on completely random vectors (crvs), see, e.g., Kallenberg (2017). The proposed theoretical framework is very broad and therefore we focus on a notable subclass, i.e., finite completely random vectors (fcrvs), which prescribe that the number of traits in the population is finite but random. We provide a comprehensive theoretical analysis of partially exchangeable trait allocations under crvs and fcrvs, including: (i) the marginal distribution of a sample and (ii) posterior representations. To illustrate the applicability of our framework, we discuss two examples involving binary traits (i.e. features) and Poisson counts. In these key special cases, the analytical derivations are extremely tractable.
@@
To illustrate the practical relevance of our methodological framework, we analyze the criminal network dataset of the ‘Ndrangheta, previously examined in Legramanti et al. (2022); Lu et al. (2025). The data were collected during Operazione Infinito (Calderoni et al., 2017), a large-scale law enforcement initiative aimed at dismantling the core branch of the ‘Ndrangheta Mafia in the Milan area. The dataset records multivariate binary outcomes describing the attendance of known affiliates (subjects) at a series of meetings (binary traits, i.e., features). In addition, affiliates can be grouped according to their locali affiliation, as documented in juridical records. This partition naturally suggests a partially exchangeable (or known-groups) framework, where inference can be carried out in closed form by leveraging our theoretical results. A distinctive feature of the proposed trait allocation model, in contrast with classical approaches to multivariate count data, is that the number of traits, that is, the columns of the data matrix, is itself a random variable. In our motivating application, this means that some meetings may remain unobserved because they were not detected by law enforcement. Our methodology explicitly accounts for this possibility while also enabling the estimation of the number of unseen traits.
@@
Unfortunately, the locali partition, while highly informative, is insufficient to fully capture the complexity of the relationships among subjects (Legramanti et al., 2022; Lu et al., 2025). To address this, we extend the partially exchangeable model by allowing the partition structure itself to be inferred from the data, thereby leading to a clustering problem. In this setting, the groups are not predetermined but are learned from the data; for this reason, we refer to it as the unknown-groups framework. The second main contribution of this paper is thus a novel mixture model designed to cluster trait allocations. Inference is enabled by the distributional results established for the simpler known-groups setting, leading to an efficient Gibbs sampling strategy. Importantly, by treating the partition as a random quantity, the proposed model generalizes Bayesian nonparametric latent class models, such as Dunson and Xing (2009). A key distinction, however, is that the total number of traits is unknown and must be estimated from the data. We formally demonstrate that ignoring this aspect inevitably results in overclustering.
@@ 2 Background on exchangeable trait allocation models
We begin by reviewing exchangeable trait allocation models (Campbell et al., 2018), in which data for each of the $n$ subjects are conditionally independent and identically distributed (iid) draws from a common and simple distribution. These highly tractable models serve as building blocks for the more flexible approaches introduced later in Sections 3-4.
@@
The exchangeable setting, especially when traits are binary, has been thoroughly investigated, e.g., in James (2017); Campbell et al. (2018), and recently by Ghilotti et al. (2025); Beraha et al. (2025) in the binary case. In the following, we present two relevant examples for the distribution $P(\cdot;\theta)$.
@@ 3.1 Model specification and completely random vectors
In equations (5)–(6), we specified a prior distribution $\mathcal{Q}_{d}$ for the vector of random measures $\tilde{\bm{\mu}}$ in (4) as a finite collection of random variables. This specification is both transparent and constructive, but it does not make explicit the connection with infinite-dimensional trait models (e.g. James, 2017; Shen et al., 2024). In Section S1 of the Supplementary Material, we prove that $\tilde{\bm{\mu}}$ is a finite completely random vector, which is a special case of the broader class of completely random vectors; see Kallenberg (2017) for a comprehensive treatment and Section S1 of the Supplementary Material for a concise overview. This result crucially relies on the Poisson specification for $N$. More precisely, $\tilde{\bm{\mu}}$ can be interpreted as an fcrv with parameters $H^{(d)}$, $\lambda$, and $P_{0}$, that is, a crv whose Lévy intensity takes the form $H^{(d)}(\mathrm{d}\theta_{1}\,\ldots\,\mathrm{d}\theta_{d})\cdot\lambda P_{0}(\mathrm{d}x)$, and we write $\tilde{\bm{\mu}}\sim\textsc{fcrv}(H^{(d)},\lambda,P_{0})$. This connection to the theory of completely random vectors is helpful, since the results presented in Section 3.2 follow as special cases of this general theory. In Section S2 of the Supplementary Material we present results for the whole class of crvs characterized by Lévy intensities of the form $\rho_{d}(\mathrm{d}\theta_{1}\,\ldots\,\mathrm{d}\theta_{d})\cdot\lambda P_{0}(\mathrm{d}x)$, where $\rho_{d}$ is a possibly infinite measure. This general framework encompasses both finite- and infinite-dimensional trait allocation models. The main practical challenge is identifying suitable, non-degenerate choices for $\rho_{d}$. Our proposal focuses on the finite case $\rho_{d}=H^{(d)}$ with $H^{(d)}$ being a probability distribution, which allows for the estimation of the total number of traits, while Shen et al. (2024) provide an alternative construction in the infinite-dimensional setting.
@@ 3.2 Distribution theory and posterior inference
A rigorous and general proof of Theorem 1, not based on the above argument, is provided in the Supplementary Material, Section S2, under general crv priors, covering the case with infinitely many traits. It is interesting to note that the probabilistic quantity in (9) appears related to the partially Exchangeable Partition Probability Function (peppf) discussed in Franzolini et al. (2025) for partially exchangeable species sampling models; however, here we consider the distribution of a random trait allocation rather than a random partition. By analogy, we refer to equations (8) and (9) as a partially Exchangeable Trait Probability Functions (petpf), with the former being conditional on $N$ and the latter being a Poisson mixture of (8) over $N$.
@@
where $(a)_{n}=a(a+1)\cdots(a+n-1)$ is the ascending factorial with $a>0$. When $d=1$, this corresponds to the efpf of the Poisson mixture of Beta-Bernoulli (Ghilotti et al., 2025).
@@ 4.1 A mixture model for trait allocations
Note that the number of groups $d$ is itself a random variable. The main difference, compared to Section 3, is that we now specify a prior for the partition $\bm{C}$. Among the various options, we consider the prior induced by a Pitman–Yor process (Perman et al., 1992; Pitman and Yor, 1997), which stands for its analytical tractability and whose density is
@@
where $\gamma>-\sigma$, $\sigma\in[0,1)$. Expression (14) is referred to as exchangeable partition probability function (Pitman, 1996). When $\sigma=0$, the Pitman–Yor process reduces to the partition law of a Dirichlet process (Ferguson, 1973). For a general characterization of exchangeable Gibbs-type random partitions, including the Pitman–Yor as a special case, see De Blasi et al. (2015). Recent developments on random partitions with finitely many clusters are discussed in Lijoi et al. (2020); Argiento and De Iorio (2022); Lijoi et al. (2024); Colombi et al. (2025).
@@
As mentioned in the introduction, the Bayesian model defined in equation (16) is related to a classical statistical approach for modeling multivariate discrete data, known as latent class analysis (Lazarsfeld and Henry, 1968; Goodman, 1974; Hagenaars and McCutcheon, 2002). The more recent work of Dunson and Xing (2009) provides a Bayesian nonparametric extension of this framework, making it even more closely related to ours. However, all these approaches assume that $N$ is a known constant, meaning that all traits are known in advance, including those equal to zero for all subjects. While this may be a reasonable assumption in many contexts, such as when the number of variables is fixed, it may be untenable in others. Ignoring unseen traits can skew the analysis and directly affect clustering, as we now show.
@@ 5 Overview of simulation studies
The simulation studies serve two main purposes. The first, detailed in Section S4.1 of the Supplementary Material, provides empirical evidence for the discrepancy highlighted in Proposition 1, which can lead to substantial inferential differences. To illustrate, we consider a simulated binary-outcome example comparing our proposed unknown-groups model (16) with a latent class model fixing $N=k$, thereby excluding the possibility of unseen traits as in Dunson and Xing (2009). In line with Proposition 1, our model yields fewer clusters than this naïve specification, with a substantial difference in our comparison. The second goal, discussed in Section S4.2 of the Supplementary Material, is to assess the performance of the unknown-groups model introduced in Section 4. Motivated by criminal network analysis, we simulate data reflecting realistic network structures. In such contexts, the focus is on the weighted adjacency matrix, which records the number of meetings co-attended by pairs of affiliates (Legramanti et al., 2022; Lu et al., 2025). This matrix summarizes the raw data, originally collected as multivariate binary observations, where each affiliate is associated with the list of meetings they attended. Our framework induces a probabilistic structure on it. To evaluate model flexibility, we consider two simulated scenarios, each producing distinct adjacency matrix properties resembling those in Lu et al. (2025). Beyond inference on the adjacency matrix, a central goal is estimating the number of meetings that went undetected by investigators. We benchmark our model-based estimates against the negative binomial mixture of bbs from Ghilotti et al. (2025), a natural and simple competitor that does not take the group structure into account. Across both scenarios, the proposed model accurately recovers the underlying structure: the posterior mean of the adjacency matrix closely matches the simulated one, and the estimates of undetected meetings are reliable. In contrast, the competitor fails to capture the number of unseen features in the first scenario and the structure of the connectivity patterns in the second scenario. Its poor fit stems from its enforced homogeneity across groups, which yields a posterior expectation of the adjacency matrix assigning identical values to all entries.
@@ 6 Analysis of the Infinito network
In this section, we analyze the Infinito ‘Ndrangheta network (Legramanti et al., 2022; Lu et al., 2025). These data stem from Operazione Infinito (Calderoni et al., 2017), a six-year-long, large-scale law enforcement operation aimed at monitoring and dismantling the core structure of La Lombardia, the highly pervasive branch of the ‘Ndrangheta Mafia operating in the Milan area. The raw data collected during the investigation are publicly available at https://sites.google.com/site/ucinetsoftware/datasets/covert-networks in the form of multivariate binary outcome observations, making them well-suited for analysis using the binary traits specification of our proposed model.
@@
The attendance of the $n$ affiliates at the $k$ distinct meetings can be represented by the $n\times k$ binary matrix $\bm{A}$, where each entry $A_{i\ell}\in\{0,1\}$ indicates whether individual $i$ attended meeting $\ell$. A graphical representation of the observed weighted adjacency matrix $\bm{W}=\bm{A}\bm{A}^{T}$, whose generic entry $W_{ii^{\prime}}$ represents the number of meetings jointly attended by affiliates $i$ and $i^{\prime}$, is shown in Figure 3(a). Legramanti et al. (2022); Lu et al. (2025) directly modelled the matrix $\bm{W}$ without considering the information of the binary traits $A_{i\ell}$. Instead, we model $A_{i\ell}$, which will induce a model for $\bm{W}$ as a by-product, therefore making use of all the available information. We focus on the $d=5$ most populated locali, with sizes $n_{1}=16,n_{2}=14,n_{3}=23,n_{4}=15,n_{5}=16$, of the Infinito dataset, for a total of $n=84$ affiliates and $k=44$ recorded meetings among them. The affiliates’ locale membership provides a natural partition into five groups, making it reasonable to begin the analysis by leveraging this a priori known clustering. This motivates the known-groups framework described in Section 3. In particular, we consider model (4) for binary traits, where $H(\cdot;\psi)$ is the beta distribution with parameters $(-\alpha,\alpha+\beta)$, with $\psi=(\alpha,\beta)$. Prior distributions for the model parameters $\lambda$, $\alpha$, and $\alpha+\beta$ are specified as follows. The parameter $\lambda$, governing the Poisson-distributed total number of meetings, is assigned a gamma prior with parameters $(\alpha_{\lambda},\beta_{\lambda})$, as detailed in Section 3.3. The hyperparameters $(\alpha_{\lambda},\beta_{\lambda})$ are chosen so that the prior expected value of $N$ equals $\hat{N}=1.5k$, where $k=44$ denotes the observed number of meetings in the dataset, and the prior variance of $N$ is set to $10\hat{N}$. For the parameters $(a,b)=(-\alpha,\alpha+\beta)$ of the beta distribution $H$, independent gamma priors are assumed with hyperparameters $(\alpha_{a},\beta_{a})$ and $(\alpha_{b},\beta_{b})$, respectively. Specifically, $(\alpha_{a},\beta_{a})$ are set to induce a prior mean of $0.2$ for $a$ with a large variance, while $(\alpha_{b},\beta_{b})$ are chosen so that the prior mean of $b$ is $10$, also with high variance. Posterior inference relies on 10,000 iterations of a vanilla mcmc algorithm, with the first 1,000 samples discarded as burn-in and a thinning interval of 2.
@@
Figure 3(c) displays the posterior distribution of the number of clusters among affiliates. To obtain a point estimate of the clustering structure, we follow Wade and Ghahramani (2018) and select the configuration that minimizes the posterior expectation of the Variation of Information (vi) metric (Meilă, 2003). The resulting clustering consists of eight unbalanced clusters, represented by the blocks in the colored bar of Figure 3(a). The row colors indicate the affiliates’ membership to the five locali, with darker shades denoting bosses and lighter tones representing lower-ranked affiliates. As evident from Figure 3, the clustering structure learned from the data deviates from the one induced by locali membership, and the posterior expected weighted adjacency matrix in Figure 3(d) nicely aligns with the observed one. These considerations qualitatively support estimating the clustering structure from the data, rather than relying on the partition induced by the locali membership. To strengthen this argument, we provide a complementary quantitative assessment. Following standard model selection practices, we compare the waic (Watanabe-Akaike Information Criteria) (Watanabe, 2013) between the two models. Specifically, the unknown-groups model yields a waic of 7236-7236, whereas the known-groups model results in a waic of 6573-6573, strongly indicating a preference for data-driven clustering estimation.
@@
Figure 4: Posterior distributions of the number of unseen meetings in the following cases: (a) unknown-groups model with clustering learned from the data; (b) known-groups model with partition induced by locali membership; (c) negative binomial mixture of bbs (Ghilotti et al., 2025).

### v2

@@ 1 Introduction
Feature allocation models gained increasing popularity within the Bayesian nonparametric community since the definition of the Indian buffet process (ibp) by Gri06; Gri11. In these models, each subject (or observation) is characterized by the presence of a set of latent characteristics, referred to as features. A key development was due to Thi07, who determined the de Finetti measure of the ibp. These seminal contributions paved the way for a rigorous theoretical foundation of exchangeable feature allocations (Bro(13)). Relevant extensions are the three-parameter ibp (Teh(09)), scaled processes (Mas22; Camerlenghi2024), and product-form feature models defined by Bat18 and studied by Ghilotti2025. See also beraha25 for a perspective based on point processes theory and predictive characterizations for feature models. Trait allocation models (Jam(17); Camp(18)) are an important generalization of feature allocation models, where each subject’s characteristics, referred to as traits, are associated with non-negative measurements rather than binary features. Notable examples of trait allocation models are described in Bro15; Heaukulani2016; Beraha2023. Applications of feature and trait allocation models encompass several fields, including genomics (Beraha2023; Mas22), Bayesian factor analysis (Aye21), ecology (Stolf2025; Ghilotti2025), gene expression modeling (Kno11), document classification and topic modeling (zhou2016priors; Wil10), image segmentation and object recognition (Gri11; Bro15).
@@
Despite their popularity, most existing contributions on both feature and trait allocation models assume exchangeability among subjects, which implies a form of homogeneity across observations. In many applied areas, this assumption is restrictive, as data come from multiple related studies and are intrinsically heterogeneous. In such cases, the notion of partial exchangeability is more appropriate. Under this assumption, observations are divided into groups that are exchangeable within, but not across, the different groups. Partial exchangeability has a long history to handle data from dependent populations of species: refer e.g. to Nip(14); Gri(17); Cam(19)AoS; Colombi2025; franzolini2025 for specific structures, and to Quintana2022 for a complete review. Nonetheless, only a few and recent contributions are available for trait and feature allocation models. For instance, Mas(18); Beraha2023; james2024HIBP introduced hierarchical constructions to induce dependence among groups, while shen2024 proposed a bivariate beta process.
@@
This paper introduces a novel general and tractable class of Bayesian nonparametric priors suitable for modeling partially exchangeable trait allocations. Our proposal relies on completely random vectors (crvs), see, Cat(21)AoS. Also, see Kallenberg_2017. The proposed theoretical framework is very broad and therefore we focus on a notable subclass, i.e., finite completely random vectors (fcrvs), which prescribe that the number of traits in the population is finite but random. We provide a comprehensive theoretical analysis of partially exchangeable trait allocations under crvs and fcrvs, including: (i) the marginal distribution of a sample and (ii) posterior representations. To illustrate the applicability of our framework, we discuss two examples involving binary traits (i.e. features) and Poisson counts. In these key special cases, the analytical derivations are extremely tractable.
@@
To illustrate the practical relevance of our methodological framework, we analyze the criminal network dataset of the ‘Ndrangheta, previously examined in esbm_rigon; Lu2025. The data were collected during Operazione Infinito (Calderoni17), a large-scale law enforcement initiative aimed at dismantling the core branch of the ‘Ndrangheta Mafia in the Milan area. The dataset records multivariate binary outcomes describing the attendance of known affiliates (subjects) at a series of meetings (binary traits, i.e., features). In addition, affiliates can be grouped according to their locali affiliation, as documented in juridical records. This partition naturally suggests a partially exchangeable (or known-groups) framework, where inference can be carried out in closed form by leveraging our theoretical results. A distinctive feature of the proposed trait allocation model, in contrast with classical approaches to multivariate count data, is that the number of traits, that is, the columns of the data matrix, is itself a random variable. In our motivating application, this means that some meetings may remain unobserved because they were not detected by law enforcement. Our methodology explicitly accounts for this possibility while also enabling the estimation of the number of unseen traits.
@@
Unfortunately, the locali partition, while highly informative, is insufficient to fully capture the complexity of the relationships among subjects (esbm_rigon; Lu2025). To address this, we extend the partially exchangeable model by allowing the partition structure itself to be inferred from the data, thereby leading to a clustering problem. In this setting, the groups are not predetermined but are learned from the data; for this reason, we refer to it as the unknown-groups framework. The second main contribution of this paper is thus a novel mixture model designed to cluster trait allocations. Inference is enabled by the distributional results established for the simpler known-groups setting, leading to an efficient Gibbs sampling strategy. Importantly, by treating the partition as a random quantity, the proposed model generalizes Bayesian nonparametric latent class models, such as Dunson2009. A key distinction, however, is that the total number of traits is unknown and must be estimated from the data. We formally demonstrate that ignoring this aspect inevitably results in overclustering.
@@ 2 Background on exchangeable trait allocation models
We begin by reviewing exchangeable trait allocation models (Camp(18)), in which data for each of the $n$ subjects are conditionally independent and identically distributed (iid) draws from a common and simple distribution. These highly tractable models serve as building blocks for the more flexible approaches introduced later in Sections 3-4.
@@
The exchangeable setting, especially when traits are binary, has been thoroughly investigated, e.g., in Jam(17); Camp(18), and recently by Ghilotti2025; beraha25 in the binary case. In the following, we present two relevant examples for the distribution $P(\cdot;\theta)$.
@@ 3.1 Model specification and completely random vectors
In equations (5)–(6), we specified a prior distribution $\mathcal{Q}_{d}$ for the vector of random measures $\tilde{\bm{\mu}}$ in (4) as a finite collection of random variables. This specification is both transparent and constructive, but it does not make explicit the connection with infinite-dimensional trait models (e.g. Jam(17); shen2024). In Section S1 of the Supplementary Material, we prove that $\tilde{\bm{\mu}}$ is a finite completely random vector, which is a special case of the broader class of completely random vectors (Cat(21)AoS); see also Kallenberg_2017 for a comprehensive treatment and Section S1 of the Supplementary Material for a concise overview. This result crucially relies on the Poisson specification for $N$. More precisely, $\tilde{\bm{\mu}}$ can be interpreted as an fcrv with parameters $H^{(d)}$, $\lambda$, and $P_{0}$, that is, a crv whose Lévy intensity takes the form $H^{(d)}(\mathrm{d}\theta_{1}\,\ldots\,\mathrm{d}\theta_{d})\cdot\lambda P_{0}(\mathrm{d}x)$, and we write $\tilde{\bm{\mu}}\sim\textsc{fcrv}(H^{(d)},\lambda,P_{0})$. This connection to the theory of completely random vectors is helpful, since the results presented in Section 3.2 follow as special cases of this general theory. In Section S2 of the Supplementary Material we present results for the whole class of crvs characterized by Lévy intensities of the form $\rho_{d}(\mathrm{d}\theta_{1}\,\ldots\,\mathrm{d}\theta_{d})\cdot\lambda P_{0}(\mathrm{d}x)$, where $\rho_{d}$ is a possibly infinite measure. This general framework encompasses both finite- and infinite-dimensional trait allocation models. The main practical challenge is identifying suitable, non-degenerate choices for $\rho_{d}$. Our proposal focuses on the finite case $\rho_{d}=H^{(d)}$ with $H^{(d)}$ being a probability distribution, which allows for the estimation of the total number of traits, while shen2024 provide an alternative construction in the infinite-dimensional setting.
@@ 3.2 Distribution theory and posterior inference
A rigorous and general proof of Theorem 1, not based on the above argument, is provided in the Supplementary Material, Section S2, under general crv priors, covering the case with infinitely many traits. It is interesting to note that the probabilistic quantity in (9) appears related to the partially Exchangeable Partition Probability Function (peppf) discussed in franzolini2025 for partially exchangeable species sampling models; however, here we consider the distribution of a random trait allocation rather than a random partition. By analogy, we refer to equations (8) and (9) as a partially Exchangeable Trait Probability Functions (petpf), with the former being conditional on $N$ and the latter being a Poisson mixture of (8) over $N$.
@@
where $(a)_{n}=a(a+1)\cdots(a+n-1)$ is the ascending factorial with $a>0$. When $d=1$, this corresponds to the efpf of the Poisson mixture of Beta-Bernoulli (Ghilotti2025).
@@ 4.1 A mixture model for trait allocations
Note that the number of groups $d$ is itself a random variable. The main difference, compared to Section 3, is that we now specify a prior for the partition $\bm{C}$. Among the various options, we consider the prior induced by a Pitman–Yor process (Per(92); Pit(97)), which stands for its analytical tractability and whose density is
@@
where $\gamma>-\sigma$, $\sigma\in[0,1)$. Expression (14) is referred to as exchangeable partition probability function (Pitman1996). When $\sigma=0$, the Pitman–Yor process reduces to the partition law of a Dirichlet process (Ferguson1973). For a general characterization of exchangeable Gibbs-type random partitions, including the Pitman–Yor as a special case, see DeBlasi2015. Recent developments on random partitions with finitely many clusters are discussed in Lijoi2020; Arg(22); Lijoi2024; Colombi2025.
@@
As mentioned in the introduction, the Bayesian model defined in equation (16) is related to a classical statistical approach for modeling multivariate discrete data, known as latent class analysis (Lazarsfeld1968; Goodman1974; Hagenaars2002). The more recent work of Dunson2009 provides a Bayesian nonparametric extension of this framework, making it even more closely related to ours. However, all these approaches assume that $N$ is a known constant, meaning that all traits are known in advance, including those equal to zero for all subjects. While this may be a reasonable assumption in many contexts, such as when the number of variables is fixed, it may be untenable in others. Ignoring unseen traits can skew the analysis and directly affect clustering, as we now show.
@@ 5 Overview of simulation studies
The simulation studies serve two main purposes. The first, detailed in Section S4.1 of the Supplementary Material, provides empirical evidence for the discrepancy highlighted in Proposition 1, which can lead to substantial inferential differences. To illustrate, we consider a simulated binary-outcome example comparing our proposed unknown-groups model (16) with a latent class model fixing $N=k$, thereby excluding the possibility of unseen traits as in Dunson2009. In line with Proposition 1, our model yields fewer clusters than this naïve specification, with a substantial difference in our comparison. The second goal, discussed in Section S4.2 of the Supplementary Material, is to assess the performance of the unknown-groups model introduced in Section 4. Motivated by criminal network analysis, we simulate data reflecting realistic network structures. In such contexts, the focus is on the weighted adjacency matrix, which records the number of meetings co-attended by pairs of affiliates (esbm_rigon; Lu2025). This matrix summarizes the raw data, originally collected as multivariate binary observations, where each affiliate is associated with the list of meetings they attended. Our framework induces a probabilistic structure on it. To evaluate model flexibility, we consider two simulated scenarios, each producing distinct adjacency matrix properties resembling those in Lu2025. Beyond inference on the adjacency matrix, a central goal is estimating the number of meetings that went undetected by investigators. We benchmark our model-based estimates against the negative binomial mixture of bbs from Ghilotti2025, a natural and simple competitor that does not take the group structure into account. Across both scenarios, the proposed model accurately recovers the underlying structure: the posterior mean of the adjacency matrix closely matches the simulated one, and the estimates of undetected meetings are reliable. In contrast, the competitor fails to capture the number of unseen features in the first scenario and the structure of the connectivity patterns in the second scenario. Its poor fit stems from its enforced homogeneity across groups, which yields a posterior expectation of the adjacency matrix assigning identical values to all entries.
@@ 6 Analysis of the Infinito network
In this section, we analyze the Infinito ‘Ndrangheta network (esbm_rigon; Lu2025). These data stem from Operazione Infinito (Calderoni17), a six-year-long, large-scale law enforcement operation aimed at monitoring and dismantling the core structure of La Lombardia, the highly pervasive branch of the ‘Ndrangheta Mafia operating in the Milan area. The raw data collected during the investigation are publicly available at https://sites.google.com/site/ucinetsoftware/datasets/covert-networks in the form of multivariate binary outcome observations, making them well-suited for analysis using the binary traits specification of our proposed model.
@@
The attendance of the $n$ affiliates at the $k$ distinct meetings can be represented by the $n\times k$ binary matrix $\bm{A}$, where each entry $A_{i\ell}\in\{0,1\}$ indicates whether individual $i$ attended meeting $\ell$. A graphical representation of the observed weighted adjacency matrix $\bm{W}=\bm{A}\bm{A}^{T}$, whose generic entry $W_{ii^{\prime}}$ represents the number of meetings jointly attended by affiliates $i$ and $i^{\prime}$, is shown in Figure 3(a). esbm_rigon; Lu2025 directly modelled the matrix $\bm{W}$ without considering the information of the binary traits $A_{i\ell}$. Instead, we model $A_{i\ell}$, which will induce a model for $\bm{W}$ as a by-product, therefore making use of all the available information. We focus on the $d=5$ most populated locali, with sizes $n_{1}=16,n_{2}=14,n_{3}=23,n_{4}=15,n_{5}=16$, of the Infinito dataset, for a total of $n=84$ affiliates and $k=44$ recorded meetings among them. The affiliates’ locale membership provides a natural partition into five groups, making it reasonable to begin the analysis by leveraging this a priori known clustering. This motivates the known-groups framework described in Section 3. In particular, we consider model (4) for binary traits, where $H(\cdot;\psi)$ is the beta distribution with parameters $(-\alpha,\alpha+\beta)$, with $\psi=(\alpha,\beta)$. Prior distributions for the model parameters $\lambda$, $\alpha$, and $\alpha+\beta$ are specified as follows. The parameter $\lambda$, governing the Poisson-distributed total number of meetings, is assigned a gamma prior with parameters $(\alpha_{\lambda},\beta_{\lambda})$, as detailed in Section 3.3. The hyperparameters $(\alpha_{\lambda},\beta_{\lambda})$ are chosen so that the prior expected value of $N$ equals $\hat{N}=1.5k$, where $k=44$ denotes the observed number of meetings in the dataset, and the prior variance of $N$ is set to $10\hat{N}$. For the parameters $(a,b)=(-\alpha,\alpha+\beta)$ of the beta distribution $H$, independent gamma priors are assumed with hyperparameters $(\alpha_{a},\beta_{a})$ and $(\alpha_{b},\beta_{b})$, respectively. Specifically, $(\alpha_{a},\beta_{a})$ are set to induce a prior mean of $0.2$ for $a$ with a large variance, while $(\alpha_{b},\beta_{b})$ are chosen so that the prior mean of $b$ is $10$, also with high variance. Posterior inference relies on 10,000 iterations of a vanilla mcmc algorithm, with the first 1,000 samples discarded as burn-in and a thinning interval of 2.
@@
Figure 3(c) displays the posterior distribution of the number of clusters among affiliates. To obtain a point estimate of the clustering structure, we follow Wad(18) and select the configuration that minimizes the posterior expectation of the Variation of Information (vi) metric (Mei(03)). The resulting clustering consists of eight unbalanced clusters, represented by the blocks in the colored bar of Figure 3(a). The row colors indicate the affiliates’ membership to the five locali, with darker shades denoting bosses and lighter tones representing lower-ranked affiliates. As evident from Figure 3, the clustering structure learned from the data deviates from the one induced by locali membership, and the posterior expected weighted adjacency matrix in Figure 3(d) nicely aligns with the observed one. These considerations qualitatively support estimating the clustering structure from the data, rather than relying on the partition induced by the locali membership. To strengthen this argument, we provide a complementary quantitative assessment. Following standard model selection practices, we compare the waic (Watanabe-Akaike Information Criteria) (Watanabe13) between the two models. Specifically, the unknown-groups model yields a waic of 7236-7236, whereas the known-groups model results in a waic of 6573-6573, strongly indicating a preference for data-driven clustering estimation.
@@
Figure 4: Posterior distributions of the number of unseen meetings in the following cases: (a) unknown-groups model with clustering learned from the data; (b) known-groups model with partition induced by locali membership; (c) negative binomial mixture of bbs (Ghilotti2025).